# Qingchuan Yang 1,  Simon Mahns  [^2] footnotemark: 2,  Sida Li 322footnotemark: 2,  Anri Gu 322footnotemark: 2,  Jibang Wu 4,  Haifeng Xu  [^3] footnotemark: 3

### v1

@@ C.7.1 Knowledge Internalization Events
We sample 100 events from Kalshi (Inc., 2025), with market close time before October 2023 (i.e., before all model knowledge cutoff dates. Note that despite that many popular events on Kalshi are Sports events as of August 2025, sports betting was not legal on Kalshi until 2024 (Wilmot, 2025). The sampled past events span the categories available on the platform during that period and exhibit differing levels of temporal granularity. Some events target a specific timestamp or date (e.g., NASDAQ price on August 20, 2023), while others are coarser, period-level questions without a single focal timestamp (e.g., Will WTI crude oil prices decrease in Q2 2023?).

### v2

@@ C.7.1 Knowledge Internalization Events
We sample 100 events from Kalshi (kalshi), with market close time before October 2023 (i.e., before all model knowledge cutoff dates. Note that despite that many popular events on Kalshi are Sports events as of August 2025, sports betting was not legal on Kalshi until 2024 (wilmot2025kalshi). The sampled past events span the categories available on the platform during that period and exhibit differing levels of temporal granularity. Some events target a specific timestamp or date (e.g., NASDAQ price on August 20, 2023), while others are coarser, period-level questions without a single focal timestamp (e.g., Will WTI crude oil prices decrease in Q2 2023?).